Paul is 3 times as old as Laura, while Sherry is 6 years older than Laura. If Paul and Sherry are twins, what is Paul's present age?
Answer by mananth(16949)

Paul is 3 times as old as Laura, while Sherry is 6 years older than Laura. If Paul and Sherry are twins, what is Paul's present age?
let laura be x years
sherry = x+6
Paul = 3x
3x= x+6
3x-x=6
2x=6
x=3
Paul is 3x
Paul is 9 years